# Assessing treatment efficacy for interval-censored endpoints using multistate semi-Markov models fit to multiple data streams

### v2

@@ A.3 MCEM effective sample size
From Liu and Liu (2001), Section 2.5.3, the effective sample size associated with the importance sampling estimate (S4) for person $i$ is, in terms of the self-normalized weights $\widetilde{\nu}_{i,m}=\nu_{i,m}/\sum_{j=1}^{M_{i}}\nu_{i,j}$, $m=1,\dots,M_{i}$,
$\textsf{ESS}_{i}$ $=\frac{M_{i}}{1+M_{i}^{2}Var_{h^{\prime}_{i}}(\widetilde{\nu}_{i})}$
$=\frac{M_{i}}{1+M_{i}^{2}[E_{h^{\prime}_{i}}(\widetilde{\nu}_{i}^{2})-E_{h^{\prime}_{i}}^{2}(\widetilde{\nu}_{i})]}$
$=\frac{M_{i}}{1+M_{i}^{2}[E_{h^{\prime}_{i}}(\widetilde{\nu}_{i}^{2})-1/M_{i}^{2}]}$
$=\frac{1}{M_{i}E_{h^{\prime}_{i}}(\widetilde{\nu}_{i}^{2})}$
$\approx\frac{1}{\sum_{j=1}^{M_{i}}\widetilde{\nu}_{i,m}^{2}}=M^{\star}_{i},$
where the third equality holds because $E_{h^{\prime}}(\widetilde{\nu}_{i})=1/M_{i}$ by construction. If the proposal distribution for the sample paths, $h^{\prime}_{i}$ is similar to the target $f^{\prime}_{i}$, then $M^{\star}_{i}$ will be close to the sample size $M_{i}$, indicating that the importance sampling sampling procedure is almost as efficient as standard Monte Carlo; while if there is a large discrepancy between $h^{\prime}_{i}$ and $f^{\prime}_{i}$, then $M^{\star}_{i}$ may be much smaller than $M_{i}$, indicating that a large number of sample paths may be necessary to obtain a stable estimate of the expected log-likelihood (S4).
@@ A.4 Ascent MCEM decision rules
Let $\widehat{\sigma}_{ASE}/{M_{Eff}}$ be the estimated asymptotic standard error (ASE) for the change in the Q-function with a target effective sample size of $M_{Eff}$, $\Delta Q_{M_{Eff}}\left(\widehat{\boldsymbol{\theta}}^{(t)},\widehat{\boldsymbol{\theta}}^{(t-1)}\right)=Q_{M_{Eff}}\left(\widehat{\boldsymbol{\theta}}^{(t)},\widehat{\boldsymbol{\theta}}^{(t-1)}\right)-Q_{M_{Eff}}\left(\widehat{\boldsymbol{\theta}}^{(t-1)},\widehat{\boldsymbol{\theta}}^{(t-1)}\right)$, the formula for which is given in equation (14) of Caffo and others (2005).
@@ A.4.1 MCEM ascent determination
$\Delta Q_{M_{Eff}}\left(\widehat{\boldsymbol{\theta}}^{(t)},\widehat{\boldsymbol{\theta}}^{(t-1)}\right)-z_{\alpha}\widehat{\sigma}_{ASE}/\sqrt{M_{Eff}}>0.$ (S5)
If the lower bound of the asymptotic confidence interval on $\Delta Q_{M_{Eff}}\left(\widehat{\boldsymbol{\theta}}^{(t)},\widehat{\boldsymbol{\theta}}^{(t-1)}\right)$ is negative, we increase the target ESS for each individual by a factor $\kappa>1$. The MCEM E- and M- steps are then repeated with the augmented pool of samples. We iterate between augmenting the pool of samples and the E- and M- steps until (S5) is satisfied, at which point we accept the maximizer.
@@ A.4.2 MCEM stopping rule
$\Delta Q_{M_{Eff}}\left(\widehat{\boldsymbol{\theta}}^{(t)},\widehat{\boldsymbol{\theta}}^{(t-1)}\right)+z_{\gamma}\widehat{\sigma}_{ASE}/\sqrt{M_{Eff}}<C,$ (S6)
i.e., if the upper bound of the confidence interval for $\Delta Q_{M_{Eff}}\left(\widehat{\boldsymbol{\theta}}^{(t)},\widehat{\boldsymbol{\theta}}^{(t-1)}\right)$ excludes $C.$
@@ A.5 Observed Fisher information
where $\nabla_{\boldsymbol{\theta}}\log f(\mathbf{Z};\boldsymbol{\theta})$ and $\nabla^{2}_{\boldsymbol{\theta}}\log f(\mathbf{Z};\boldsymbol{\theta})$ are the gradient and hessian of the log-likelihood under the semi-Markov model and the expectations are taken over the distribution of $\mathbf{Z}$. Given $M$ samples from the Markov surrogate drawn conditionally on the data, $Z^{(1)}_{i},\dots,Z^{(M_{i})}_{i}\sim h^{\prime}_{i}(Z\mid\mathbf{Y},\boldsymbol{\theta}^{\prime})$ with self-normalized importance weights, $\widetilde{\nu}_{i,1},\dots,\widetilde{\nu}_{i,M_{i}}$, from the final iteration of the MCEM algorithm, the observed fisher information can be approximated at the MLE, $\widehat{\boldsymbol{\theta}}$, via
$\widehat{I}(\widehat{\boldsymbol{\theta}}\mid\mathbf{Y})$ $=\sum_{i=1}^{P}M_{i}\sum_{m=1}^{M_{i}}\widetilde{\nu}_{i,m}\left[-\nabla^{2}_{\widehat{\boldsymbol{\theta}}}\log f\left(Z^{(m)}_{i};\widehat{\boldsymbol{\theta}}\right)-\left(\nabla_{\widehat{\boldsymbol{\theta}}}\log f\left(Z^{(m)}_{i};\widehat{\boldsymbol{\theta}}\right)\right)\left(\nabla_{\widehat{\boldsymbol{\theta}}}\log f\left(Z^{(m)}_{i};\widehat{\boldsymbol{\theta}}\right)\right)^{T}\right]$
$\hskip 20.00003pt+M_{i}^{2}\sum_{m=1}^{M_{i}}\sum_{n=1}^{M_{i}}\widetilde{\nu}_{i,m}\widetilde{\nu}_{i,n}\left[\nabla_{\widehat{\boldsymbol{\theta}}}\log f\left(Z^{(m)}_{i};\widehat{\boldsymbol{\theta}}\right)\right]\left[\nabla_{\widehat{\boldsymbol{\theta}}}\log f\left(Z^{(n)}_{i};\widehat{\boldsymbol{\theta}}\right)\right]^{T},$ (S8)
@@ A.6 Marginal likelihood estimation
To simplify notation, we suppress the dependence on the covariate $X_{i}$. Under a semi-Markov model, the marginal density of the observed data for participant $i=1,\dots,n$ conditional on $\widehat{\boldsymbol{\theta}}$ is
@@
$\widehat{f}_{i}\left(\mathbf{Y}_{i}\mid\boldsymbol{\theta}^{(t)}\right)=\frac{r_{i}\left(\mathbf{Y}_{i}\mid\boldsymbol{\theta}^{\prime}\right)}{M_{i}}\sum_{m=1}^{M_{i}}\nu_{i,m}^{(t)}=r_{i}\left(\mathbf{Y}_{i}\mid\boldsymbol{\theta}^{\prime}\right)\bar{\nu}_{i}^{(t)}$ (S10)
@@
$\widehat{f}\left(\mathbf{Y}\mid\boldsymbol{\theta}^{(t)}\right)=\prod_{i=1}^{P}\widehat{f}_{i}\left(\mathbf{Y}_{i}\mid\boldsymbol{\theta}^{(t)}\right)=r(\mathbf{Y}\mid\boldsymbol{\theta}^{\prime})\prod_{i=1}^{P}\bar{\nu}_{i}^{(t)},$ (S11)
@@
$\textsf{Var}\left(\widehat{f}_{i}\left(\mathbf{Y}_{i}\mid\boldsymbol{\theta}^{(t)}\right)\right)=(r_{i}(\mathbf{Y}_{i}\mid\boldsymbol{\theta}^{\prime}))^{2}\textsf{Var}\left(\bar{\nu}_{i}^{(t)}\right),$
@@
$\widehat{\textsf{Var}}\left(\widehat{f}_{i}\left(\mathbf{Y}_{i}\mid\boldsymbol{\theta}^{(t)}\right)\right)=(r_{i}(\mathbf{Y}_{i}\mid\boldsymbol{\theta}^{\prime}))^{2}\widehat{\textsf{Var}}\left(\bar{\nu}_{i}^{(t)}\right),$
@@
$\widehat{\textsf{Var}}\left(\bar{\nu}_{i}^{(t)}\right)=\frac{1}{M_{i}}\frac{1}{M_{i}-1}\sum_{m=1}^{M_{i}}\left(\nu_{i,m}^{(t)}-\bar{\nu}_{i}^{(t)}\right)^{2}.$
@@
$\log\widehat{f}\left(\mathbf{Y}\mid\boldsymbol{\theta}^{(t^{\star})}\right)=\log r(\mathbf{Y}\mid\boldsymbol{\theta}^{\prime})+\sum_{i=1}^{P}\log\bar{\nu}_{i}^{(t^{\star})}.$ (S12)
@@
$\widehat{\textsf{Var}}\left(\log\widehat{f}\left(\mathbf{Y}\mid\boldsymbol{\theta}^{(t^{\star})}\right)\right)=\sum_{i=1}^{P}\frac{\textsf{Var}\left(\bar{\nu}_{i}^{(t^{\star})}\right)}{\left(\bar{\nu}_{i}^{(t^{\star})}\right)^{2}}.$ (S13)
@@ B.1 Simulation model
All three transitions had Weibull baseline intensities with parameters given in Table S5. The true values for the quantities of interest estimated in the simulation in Section 3.1 are given in Table S6
@@ C.4 Additional results
Table S13: Bias, coverage, and 95% confidence interval width of estimates of key quantities obtained by fitting the fully parametric 5 state model shown in 1(c) and S11 to panel data where PCR and serology were assessed at study day 28. RM TI is the restricted mean time to infection, RM PCR+ is the restricted mean duration of PCR positivity, RR is relative risk, and PE is protective efficacy, calculated as $PE=1-RR$.
@@ D.3 Additional results
Table S18: Additional results from the analysis of REGEN-2069 data. We abbreviate as follows: Pr = Probability, RMIFT = restricted mean infection free time, RM PCR+ = restricted mean duration of PCR+, PE = protective efficacy, RR = relative risk, N+ = seropositive. RM TI is the arithmetic mean over individuals of the the predicted days to infection or maximum follow-up for each individual, whichever is less. RM PCR+ is the arithmetic mean days during which individuals are detectable by PCR or the maximum follow-up for each individual, whichever is less. RR is the ratio of probabilities of the specified event, mAb divided by placebo. PE is 1 - RR, i.e., the reduction in relative risk.

### v3

@@ A.3 MCEM effective sample size
From Liu and Liu (2001), Section 2.5.3, the effective sample size associated with the importance sampling estimate (S4) for person $i$ is, in terms of the self-normalized weights $\tilde{\nu}_{i,m}=\nu_{i,m}/\sum_{j=1}^{M_{i}}\nu_{i,j}$, $m=1,\dots,M_{i}$,
$\textsf{ESS}_{i}$ $=\frac{M_{i}}{1+M_{i}^{2}Var_{h^{\prime}_{i}}(\tilde{\nu}_{i})}$
$=\frac{M_{i}}{1+M_{i}^{2}[E_{h^{\prime}_{i}}(\tilde{\nu}_{i}^{2})-E_{h^{\prime}_{i}}^{2}(\tilde{\nu}_{i})]}$
$=\frac{M_{i}}{1+M_{i}^{2}[E_{h^{\prime}_{i}}(\tilde{\nu}_{i}^{2})-1/M_{i}^{2}]}$
$=\frac{1}{M_{i}E_{h^{\prime}_{i}}(\tilde{\nu}_{i}^{2})}$
$\approx\frac{1}{\sum_{j=1}^{M_{i}}\tilde{\nu}_{i,m}^{2}}=M^{\star}_{i},$
where the third equality holds because $E_{h^{\prime}}(\tilde{\nu}_{i})=1/M_{i}$ by construction. If the proposal distribution for the sample paths, $h^{\prime}_{i}$ is similar to the target $f^{\prime}_{i}$, then $M^{\star}_{i}$ will be close to the sample size $M_{i}$, indicating that the importance sampling procedure is almost as efficient as standard Monte Carlo; while if there is a large discrepancy between $h^{\prime}_{i}$ and $f^{\prime}_{i}$, then $M^{\star}_{i}$ may be much smaller than $M_{i}$, indicating that a large number of sample paths may be necessary to obtain a stable estimate of the expected log-likelihood (S4).
@@ A.4 Ascent MCEM decision rules
Let $\hat{\sigma}_{ASE}/{M_{Eff}}$ be the estimated asymptotic standard error (ASE) for the change in the Q-function with a target effective sample size of $M_{Eff}$, $\Delta Q_{M_{Eff}}\left(\hat{\boldsymbol{\theta}}^{(t)},\hat{\boldsymbol{\theta}}^{(t-1)}\right)=Q_{M_{Eff}}\left(\hat{\boldsymbol{\theta}}^{(t)},\hat{\boldsymbol{\theta}}^{(t-1)}\right)-Q_{M_{Eff}}\left(\hat{\boldsymbol{\theta}}^{(t-1)},\hat{\boldsymbol{\theta}}^{(t-1)}\right)$, the formula for which is given in equation (14) of Caffo and others (2005).
@@ A.4.1 MCEM ascent determination
$\Delta Q_{M_{Eff}}\left(\hat{\boldsymbol{\theta}}^{(t)},\hat{\boldsymbol{\theta}}^{(t-1)}\right)-z_{\alpha}\hat{\sigma}_{ASE}/\sqrt{M_{Eff}}>0.$ (S5)
If the lower bound of the asymptotic confidence interval on $\Delta Q_{M_{Eff}}\left(\hat{\boldsymbol{\theta}}^{(t)},\hat{\boldsymbol{\theta}}^{(t-1)}\right)$ is negative, we increase the target ESS for each individual by a factor $\kappa>1$. The MCEM E- and M- steps are then repeated with the augmented pool of samples. We iterate between augmenting the pool of samples and the E- and M- steps until (S5) is satisfied, at which point we accept the maximizer.
@@ A.4.2 MCEM stopping rule
$\Delta Q_{M_{Eff}}\left(\hat{\boldsymbol{\theta}}^{(t)},\hat{\boldsymbol{\theta}}^{(t-1)}\right)+z_{\gamma}\hat{\sigma}_{ASE}/\sqrt{M_{Eff}}<C,$ (S6)
i.e., if the upper bound of the confidence interval for $\Delta Q_{M_{Eff}}\left(\hat{\boldsymbol{\theta}}^{(t)},\hat{\boldsymbol{\theta}}^{(t-1)}\right)$ excludes $C.$
@@ A.5 Observed Fisher information
where $\nabla_{\boldsymbol{\theta}}\log f(\mathbf{Z};\boldsymbol{\theta})$ and $\nabla^{2}_{\boldsymbol{\theta}}\log f(\mathbf{Z};\boldsymbol{\theta})$ are the gradient and hessian of the log-likelihood under the semi-Markov model and the expectations are taken over the distribution of $\mathbf{Z}$. Given $M$ samples from the Markov surrogate drawn conditionally on the data, $Z^{(1)}_{i},\dots,Z^{(M_{i})}_{i}\sim h^{\prime}_{i}(Z\mid\mathbf{Y},\boldsymbol{\theta}^{\prime})$ with self-normalized importance weights, $\tilde{\nu}_{i,1},\dots,\tilde{\nu}_{i,M_{i}}$, from the final iteration of the MCEM algorithm, the observed fisher information can be approximated at the MLE, $\hat{\boldsymbol{\theta}}$, via
$\hat{I}(\hat{\boldsymbol{\theta}}\mid\mathbf{Y})$ $=\sum_{i=1}^{P}M_{i}\sum_{m=1}^{M_{i}}\tilde{\nu}_{i,m}\left[-\nabla^{2}_{\hat{\boldsymbol{\theta}}}\log f\left(Z^{(m)}_{i};\hat{\boldsymbol{\theta}}\right)-\left(\nabla_{\hat{\boldsymbol{\theta}}}\log f\left(Z^{(m)}_{i};\hat{\boldsymbol{\theta}}\right)\right)\left(\nabla_{\hat{\boldsymbol{\theta}}}\log f\left(Z^{(m)}_{i};\hat{\boldsymbol{\theta}}\right)\right)^{T}\right]$
$\hskip 20.00003pt+M_{i}^{2}\sum_{m=1}^{M_{i}}\sum_{n=1}^{M_{i}}\tilde{\nu}_{i,m}\tilde{\nu}_{i,n}\left[\nabla_{\hat{\boldsymbol{\theta}}}\log f\left(Z^{(m)}_{i};\hat{\boldsymbol{\theta}}\right)\right]\left[\nabla_{\hat{\boldsymbol{\theta}}}\log f\left(Z^{(n)}_{i};\hat{\boldsymbol{\theta}}\right)\right]^{T},$ (S8)
@@ A.6 Marginal likelihood estimation
To simplify notation, we suppress the dependence on the covariate $X_{i}$. Under a semi-Markov model, the marginal density of the observed data for participant $i=1,\dots,n$ conditional on $\hat{\boldsymbol{\theta}}$ is
@@
$\hat{f}_{i}\left(\mathbf{Y}_{i}\mid\boldsymbol{\theta}^{(t)}\right)=\frac{r_{i}\left(\mathbf{Y}_{i}\mid\boldsymbol{\theta}^{\prime}\right)}{M_{i}}\sum_{m=1}^{M_{i}}\nu_{i,m}^{(t)}=r_{i}\left(\mathbf{Y}_{i}\mid\boldsymbol{\theta}^{\prime}\right)\bar{\nu}_{i}^{(t)}$ (S10)
@@
$\hat{f}\left(\mathbf{Y}\mid\boldsymbol{\theta}^{(t)}\right)=\prod_{i=1}^{P}\hat{f}_{i}\left(\mathbf{Y}_{i}\mid\boldsymbol{\theta}^{(t)}\right)=r(\mathbf{Y}\mid\boldsymbol{\theta}^{\prime})\prod_{i=1}^{P}\bar{\nu}_{i}^{(t)},$ (S11)
@@
$\textsf{Var}\left(\hat{f}_{i}\left(\mathbf{Y}_{i}\mid\boldsymbol{\theta}^{(t)}\right)\right)=(r_{i}(\mathbf{Y}_{i}\mid\boldsymbol{\theta}^{\prime}))^{2}\textsf{Var}\left(\bar{\nu}_{i}^{(t)}\right),$
@@
$\hat{\textsf{Var}}\left(\hat{f}_{i}\left(\mathbf{Y}_{i}\mid\boldsymbol{\theta}^{(t)}\right)\right)=(r_{i}(\mathbf{Y}_{i}\mid\boldsymbol{\theta}^{\prime}))^{2}\hat{\textsf{Var}}\left(\bar{\nu}_{i}^{(t)}\right),$
@@
$\hat{\textsf{Var}}\left(\bar{\nu}_{i}^{(t)}\right)=\frac{1}{M_{i}}\frac{1}{M_{i}-1}\sum_{m=1}^{M_{i}}\left(\nu_{i,m}^{(t)}-\bar{\nu}_{i}^{(t)}\right)^{2}.$
@@
$\log\hat{f}\left(\mathbf{Y}\mid\boldsymbol{\theta}^{(t^{\star})}\right)=\log r(\mathbf{Y}\mid\boldsymbol{\theta}^{\prime})+\sum_{i=1}^{P}\log\bar{\nu}_{i}^{(t^{\star})}.$ (S12)
@@
$\hat{\textsf{Var}}\left(\log\hat{f}\left(\mathbf{Y}\mid\boldsymbol{\theta}^{(t^{\star})}\right)\right)=\sum_{i=1}^{P}\frac{\textsf{Var}\left(\bar{\nu}_{i}^{(t^{\star})}\right)}{\left(\bar{\nu}_{i}^{(t^{\star})}\right)^{2}}.$ (S13)
@@ B.1 Simulation model
All three transitions had Weibull baseline intensities with parameters given in Table S1. The true values for the quantities of interest estimated in the simulation in Section 3.1 are given in Table S2.
@@ C.4 Additional results
Table S9: Bias, coverage, and 95% confidence interval width of estimates of key quantities obtained by fitting the fully parametric 5 state model shown in 1(c) and S7 to panel data where PCR and serology were assessed at study day 28. RM TI is the restricted mean time to infection, RM PCR+ is the restricted mean duration of PCR positivity, RR is relative risk, and PE is protective efficacy, calculated as $PE=1-RR$.
@@ D.3 Additional results
Parameterization of baseline intensities Results
Weibull Weibull 78.7 -133.3 (0.95)
Exponential Exponential 0.0 16.0 (0.0)
Table S14: Model comparison results. The degree and interior knots of the B-Spline intensities are $\kappa$ and $\xi$, respectively. The quantity $\ell(\mathbf{Y})-\ell_{M}(\mathbf{Y})$ indicates the absolute gain in log-likelihood over the Markov model with exponential transition intensities (larger is better). $\Delta$AIC is computed as $-2(\ell(\mathbf{Y})-\ell_{M}(\mathbf{Y}))+2p$ (smaller is better), where $p$ is the number of model parameters. The log-likelihood of the Markov model is obtained numerically, while those of the semi-Markov model are estimated via Monte Carlo (Appendix A.6).
@@
Table S15: Additional results from the analysis of REGEN-2069 data. We abbreviate as follows: Pr = Probability, RMIFT = restricted mean infection free time, RM PCR+ = restricted mean duration of PCR+, PE = protective efficacy, RR = relative risk, N+ = seropositive. RM TI is the arithmetic mean over individuals of the predicted days to infection or maximum follow-up for each individual, whichever is less. RM PCR+ is the arithmetic mean days during which individuals are detectable by PCR or the maximum follow-up for each individual, whichever is less. RR is the ratio of probabilities of the specified event, mAb divided by placebo. PE is 1 - RR, i.e., the reduction in relative risk.